A company offers ID theft protection using leads obtained from client banks. Three employees work 40 hours a week on the leads, at a pay rate of $12 per hour per employee. Each employee identifies an average of 3,400 potential leads a week from a list of 4,600. An average of 6 percent of potential leads actually sign up for the service, paying a one-time fee of $80. Material costs are $1,500 per week, and overhead costs are $9,300 per week.

Calculate the multifactor productivity for this operation in fees generated per dollar of input. (Round your answer to 2 decimal places.)

What is the Multifactor productivity _____
Answer #1

Note: Please rate the answer (Upvote) and do ask if any doubt is there.

Answer:

  • Multifactor productivity = Output/Input
  • Output = (possible leads*no of workers*fee*Conversion percentage)
  • Possible leads =3400
  • No of employees = 3
  • Fee = $80
  • Conversion percentage = 6%
  • Output = (3400*3*80*6%) = $48960
  • Input= Labor cost + Material cost + Overhead cost
  • Labor cost = 3*40*12 = $1440
  • Material cost = $1500
  • Overhead cost = $9300
  • Input = 1440+1500+9300 = $12240
  • MFP = 48960/12240 = 4 (Answer)

Note: Please rate the answer (Upvote) and do ask if any doubt is there.